# Coherent Absorption Dynamics: The Dual Role of Off-Diagonal Couplings in Weakly Bound Nuclei

### v1

@@ Abstract
Disentangling reaction mechanisms in weakly bound nuclei remains a long-standing challenge, complicated by the common practice of treating absorption as an incoherent sum of channel contributions. Within the Continuum-Discretized Coupled-Channels (CDCC) framework, we derive a generalized coupled-channel optical theorem and show that the total absorption cross section, $\sigma_{\mathrm{A}}\propto-\langle\Psi|W|\Psi\rangle$, decomposes as $\sigma_{\mathrm{A}}=\sigma_{\mathrm{D}}+\sigma_{\mathrm{B}}+\sigma_{\mathrm{int}}$, where $\sigma_{\mathrm{int}}$ is a coherent interference term between channel components. For the systems and complex fragment–target optical potentials considered, $\sigma_{\mathrm{int}}$ is negative and comparable in magnitude to the direct absorption terms. The off-diagonal imaginary couplings play a dual role: they redistribute flux among channels and generate $\sigma_{\mathrm{int}}$, which is required for flux-balance consistency. In calculations for $d+{}^{93}\mathrm{Nb}$ and $^{6}\mathrm{Li}+{}^{59}\mathrm{Co}/{}^{208}\mathrm{Pb}$, retaining the full non-diagonal coupling matrix nearly doubles the breakup-channel absorption for the heavy target, while reducing the total absorption through $\sigma_{\mathrm{int}}$. We demonstrate that neglecting the off-diagonal imaginary couplings $W_{ij}$ ($i\neq j$) is not merely an approximation but leads to a systematically biased physical picture: the total absorption is overestimated while the breakup absorption component is severely underestimated. Experimental analyses that employ incoherent-sum models to extract direct and breakup cross sections from data will inherit this bias. The full coupling matrix is therefore essential for mechanism-resolved cross-section extraction, and we advocate that experimentalists adopt full-coupling CDCC calculations as the standard for consistent interpretation of absorption data in weakly bound systems.
@@ I Introduction
While full-coupling calculations are computationally feasible with modern codes, simplified approaches remain prevalent in recent literature. One common approximation retains off-diagonal couplings within the continuum subspace but neglects couplings between the elastic channel and the continuum ($W_{1j}=0$ for j>1j>1[11, 3, 17, 16, 18]; another neglects all off-diagonal imaginary couplings entirely, retaining only diagonal terms [7]. Both approximations effectively assume an incoherent reaction picture in which absorption probabilities add. Such an assumption obscures the cross terms that arise naturally from expanding $\langle\Psi|W|\Psi\rangle$ over channel components. When adopted in the interpretation of experimental data, where model-assisted separations of direct and breakup absorption are common, this approximation can introduce a systematic bias, most notably a systematic underestimation of breakup absorption.
In this work, we rigorously derive a generalized coupled-channel optical theorem within the CDCC framework to address these dynamics. This formalism respects flux conservation in the projected model space and enables the theoretical separation of the total absorption cross section into three physically distinct components: direct absorption ($\sigma_{\mathrm{D}}$) from the elastic channel, breakup absorption ($\sigma_{\mathrm{B}}$) from the continuum states, and an interference term ($\sigma_{\mathrm{int}}$) arising from off-diagonal imaginary couplings between the elastic channel and the continuum states. This theoretical separation allows us to go beyond simple incoherent summation and explicitly examine the microscopic composition of the absorption process based on the wave functions and coupling potentials.
@@
The paper is organized as follows. In Sec. II, we derive the generalized coupled-channel optical theorem and present the decomposition formalism. Section III presents numerical results for $d+{}^{93}\mathrm{Nb}$ and $^{6}\mathrm{Li}+{}^{59}\mathrm{Co}/{}^{208}\mathrm{Pb}$ systems. Finally, Sec. IV summarizes our findings and discusses their implications.
@@ II Theory
II.1 Derivation of the Coupled-Channel Optical Theorem
@@ IV Summary and Discussion
In this work, we have presented a rigorous derivation of the generalized optical theorem within the Continuum-Discretized Coupled-Channels (CDCC) framework and investigated the role of off-diagonal imaginary couplings in the absorption mechanism of weakly bound nuclei. By decomposing the total absorption cross section into direct absorption ($\sigma_{\mathrm{D}}$), breakup absorption ($\sigma_{\mathrm{B}}$), and the interference term ($\sigma_{\mathrm{int}}$), we have elucidated the complex quantum interference effects that govern the redistribution of reaction flux.

### v2

@@ Abstract
Disentangling reaction mechanisms in weakly bound nuclei remains a long-standing challenge, complicated by the common practice of treating absorption as an incoherent sum of channel contributions. Within the Continuum-Discretized Coupled-Channels (CDCC) framework, we apply the generalized optical theorem [Nucl. Phys. A 842, 48 (2010)] and show that the total absorption cross section, $\sigma_{\mathrm{A}}\propto-\langle\Psi|W|\Psi\rangle$, decomposes as $\sigma_{\mathrm{A}}=\sigma_{\mathrm{D}}+\sigma_{\mathrm{B}}+\sigma_{\mathrm{int}}$, where $\sigma_{\mathrm{int}}$ is a coherent interference term between channel components. For the systems and complex fragment–target optical potentials considered, $\sigma_{\mathrm{int}}$ is negative and comparable in magnitude to the direct absorption terms. The off-diagonal imaginary couplings play a dual role: they redistribute flux among channels and generate $\sigma_{\mathrm{int}}$, which is required for flux-balance consistency. In calculations for $d+{}^{93}\mathrm{Nb}$ and $^{6}\mathrm{Li}+{}^{59}\mathrm{Co}/{}^{208}\mathrm{Pb}$, retaining the full non-diagonal coupling matrix nearly doubles the breakup-channel absorption for the heavy target, while reducing the total absorption through $\sigma_{\mathrm{int}}$. We demonstrate that neglecting the off-diagonal imaginary couplings $W_{ij}$ ($i\neq j$) is not merely an approximation but leads to a systematically biased physical picture: the total absorption is overestimated while the breakup absorption component is severely underestimated. Experimental analyses that employ incoherent-sum models to extract direct and breakup cross sections from data will inherit this bias. The full coupling matrix is therefore essential for mechanism-resolved cross-section extraction, and we advocate that experimentalists adopt full-coupling CDCC calculations as the standard for consistent interpretation of absorption data in weakly bound systems.
@@ I Introduction
While full-coupling calculations are computationally feasible with modern codes, simplified approaches remain prevalent in recent literature. One common approximation retains off-diagonal couplings within the continuum subspace but neglects couplings between the elastic channel and the continuum ($W_{1j}=0$ for j>1j>1[2, 3, 16, 17, 6]; another neglects all off-diagonal imaginary couplings entirely, retaining only diagonal terms [8, 15]. Both approximations effectively assume an incoherent reaction picture in which absorption probabilities add. Such an assumption obscures the cross terms that arise naturally from expanding $\langle\Psi|W|\Psi\rangle$ over channel components. When adopted in the interpretation of experimental data, where model-assisted separations of direct and breakup absorption are common, this approximation can introduce a systematic bias, most notably a systematic underestimation of breakup absorption.
In this work, we apply the generalized optical theorem formalism [18] within the CDCC framework to address these dynamics. This formalism respects flux conservation in the projected model space and enables the theoretical separation of the total absorption cross section into three physically distinct components: direct absorption ($\sigma_{\mathrm{D}}$) from the elastic channel, breakup absorption ($\sigma_{\mathrm{B}}$) from the continuum states, and an interference term ($\sigma_{\mathrm{int}}$) arising from off-diagonal imaginary couplings between the elastic channel and the continuum states. This theoretical separation allows us to go beyond simple incoherent summation and explicitly examine the microscopic composition of the absorption process based on the wave functions and coupling potentials.
@@
The paper is organized as follows. In Sec. II, we present the theoretical formalism based on the generalized optical theorem and derive the absorption decomposition. Section III presents numerical results for $d+{}^{93}\mathrm{Nb}$ and $^{6}\mathrm{Li}+{}^{59}\mathrm{Co}/{}^{208}\mathrm{Pb}$ systems. Finally, Sec. IV summarizes our findings and discusses their implications.
@@ II Theory
II.1 Generalized Optical Theorem in Coupled Channels
@@ IV Summary and Discussion
In this work, we have applied the generalized optical theorem [18] within the Continuum-Discretized Coupled-Channels (CDCC) framework and investigated the role of off-diagonal imaginary couplings in the absorption mechanism of weakly bound nuclei. By decomposing the total absorption cross section into direct absorption ($\sigma_{\mathrm{D}}$), breakup absorption ($\sigma_{\mathrm{B}}$), and the interference term ($\sigma_{\mathrm{int}}$), we have elucidated the complex quantum interference effects that govern the redistribution of reaction flux.